OpenStudy (anonymous):

Rick and Mike are roommates and leave Gainesville on Interstate 75 at the same time to visit their girlfriends for a long weekend. Rick travels north and Mike travels south. If Mike's average speed is 10 mph faster than Rick's, find the speed of each if they are 207 miles apart in 1 hour and 30 minutes. Rick's speed is mph. __________ Mike's speed is mph. ____________

OpenStudy (anonymous):

d=rt

OpenStudy (cwrw238):

let ricks speed be x mph then mikes speed will be x + 10 mph their combined speed will be (2x + 10) mph as they are travelling in opposite directions speed = distance / time so 2x + 10 = 207 / 1.5 solve this for x and you've done it

OpenStudy (anonymous):

x= 64? @cwrw238

OpenStudy (anonymous):

rick= 64mph ?

OpenStudy (anonymous):

I need, mike's speed. Help <//3

OpenStudy (anonymous):

mike's speed was assumed to be x+10, now since x =64, therefore Mike's speed is 64 +10 + 74 mph.